OpenStudy (anonymous):

I need help with my algebra 2 assignmen!! I will fan and medal anyone who helps!!

OpenStudy (eli_moses):

let's see what i can help you with so far

OpenStudy (anonymous):

1. To open these doors, you must match the number and type of solutions for the following two fuctions in standard form. •f(x) = x2 + 6x – 16 •g(x) = x2 +6x + 1 Match the following descriptions of the solutions to each of the functions above. Hint: they each have their own match. •Two real irrationals solutions •Two real rationals solutions

OpenStudy (anonymous):

I just don't get it

OpenStudy (eli_moses):

hmm.......

OpenStudy (eli_moses):

is the first equation x squared or 2x?

OpenStudy (anonymous):

x squared

OpenStudy (eli_moses):

ok thanks

OpenStudy (anonymous):

yea no problem

OpenStudy (kohai):

I just did this assignment today, one moment

OpenStudy (anonymous):

Really?? @kohai

OpenStudy (kohai):

Yes. I'm probably not using the same program but the assignment I did today covered this

OpenStudy (anonymous):

yea i'm in florida virtual

OpenStudy (kohai):

Do you know how to calculate the discriminant?

OpenStudy (anonymous):

ummm kinda

OpenStudy (kohai):

Tell me what you know

OpenStudy (anonymous):

actually I don't know how

OpenStudy (kohai):

Alright, so, the formula is \[b^2-4ac\]

OpenStudy (kohai):

What terms would go in what places?

OpenStudy (anonymous):

well for f(x)=x^2+6x-16 it would be.. (1)^2-4(6)(-16) right?

OpenStudy (kohai):